OpenStudy (anonymous):

Solve for x, where x is a real number. 1) log x+ log(x-3) =1 2) |x-3|< 7 Could any of u guys please explain to me how to solve this two step by step, i am so rusty on my math.thanks so much

OpenStudy (anonymous):

in 1... u can apply laws of log m + log n = log mn.... then get inverse log both side to get rid of log... then you can solve for x...

OpenStudy (anonymous):

in 2... solve for x first and get its absolute value and test for inequality

OpenStudy (anonymous):

so i get logx(x-3)=1 and then what? @Orion1213

OpenStudy (anonymous):

inv log both side.... inv log LHS = inv log RHS, this will cancel log... inv log a = 10^a

OpenStudy (anonymous):

so it would be equal to log 10, because 10 is the base right? I'll have log x^2-3x = log 10 x^2-3x=10 x^2-3x-10 & then just factor

OpenStudy (anonymous):

@Orion1213 ^

OpenStudy (anonymous):

wrong format for inv log... it should be 10^[log x(x-3)] = 10^1 btw well get the same... yeah go on factoring...

OpenStudy (anonymous):

what is the factor of -10 which will give you a sum of -3?

OpenStudy (anonymous):

so after i do x^2-3x-10=0 we'll have (x-5)(x+2)=0 The answer is x=5, because the answer cant be negative right?@Orion1213

OpenStudy (anonymous):

yes you're correct.... \(\ddot\smile\)